OpenStudy (anonymous):

find all complex solutions 4x^2+2x-1=0

zepdrix (zepdrix):

Use your Quadratic Formula:\[\large\rm x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

zepdrix (zepdrix):

\[\large\rm x=\frac{-2\pm\sqrt{(2)^2-4(4)(-1)}}{2(4)}\]

zepdrix (zepdrix):

Hmm there are only real solutions to this polynomial, no complex.

OpenStudy (anonymous):

i have been working on this(solving quadratic equations with coplex roots) for hours i just cant figure it out

OpenStudy (anonymous):

Two real roots, maybe it's intended to be \[4x^2+2x+1\]